### v2

# Emergent Open-Vocabulary Semantic Segmentation from Off-the-shelf Vision-Language Models

Abstract
From image-text pairs, large-scale vision-language models (VLMs) learn to implicitly associate image regions with words, which prove effective for tasks like visual question answering. However, leveraging the learned association for open-vocabulary semantic segmentation remains a challenge. In this paper, we propose a simple, yet extremely effective, training-free technique, Plug-and-Play Open-Vocabulary Semantic Segmentation (PnP-OVSS) for this task. PnP-OVSS leverages a VLM with direct text-to-image cross-attention and an image-text matching loss. To balance between over-segmentation and under-segmentation, we introduce Salience Dropout; by iteratively dropping patches that the model is most attentive to, we are able to better resolve the entire extent of the segmentation mask. PnP-OVSS does not require any neural network training and performs hyperparameter tuning without the need for any segmentation annotations, even for a validation set. PnP-OVSS demonstrates substantial improvements over comparable baselines (+29.4% mIoU on Pascal VOC, +13.2% mIoU on Pascal Context, +14.0% mIoU on MS COCO, +2.4% mIoU on COCO Stuff) and even outperforms most baselines that conduct additional network training on top of pretrained VLMs. Our codebase is at https://github.com/letitiabanana/PnP-OVSS.
Figure 1: Qualitative Results of PnP-OVSS + BLIP. Images are from Pascal Context and COCO stuff. The right columns and bottom rows show the ground-truth (GT); the rest are our results. Note accurate results even on complex (trees) and small objects (last column).
1 Introduction
The classic task of semantic segmentation [20, 22] aims to classify pixels to their object types. Traditional supervised methods require dense pixel-level annotations and are restricted to recognizing a predefined set of objects. To relax these constraints, open-vocabulary semantic segmentation [17, 78, 75, 41, 77, 44, 48, 46, 85, 21, 8, 81] aspires to identify arbitrary object categories, whereas weakly supervised techniques [91, 36, 11, 31, 25, 9, 55, 76, 15, 53, 66] can acquire pixel-level localization capabilities from coarse supervision, e.g., image labels or boxes.
Large-scale vision-language models (VLMs) pretrained on image-text pairs [61, 86, 40, 37, 38, 79, 1] achieve unprecedented performance on multimodal tasks, such as describing arbitrary images and answering free-form, open-ended questions about them (either with [61, 37, 38] or without finetuning [1, 63, 19, 71]). These tasks apparently involve some ability to localize objects. For example, to answer the question “what objects appear on the table?”, the model would have to first localize the table in the image and identify the objects on it. Hence, it is reasonable to conjecture that the VLM network learns to perform open-vocabulary localization from image-text pretraining. However, distilling the localization capability from the VLMs remains an open challenge.
Most existing methods for open-vocabulary semantic segmentation (OVSS) from VLMs usually obtain single vector encodings for the visual and text inputs respectively [75, 43, 51, 74, 44, 52]. However, pooling every token into a single vector likely discards information about detailed positions of objects and words. We investigate the use of pretrained cross-attention layers for OVSS, which retain finer-grained correspondence between text and image patches.
Nevertheless, a naive application of cross-attention between the class name and the image patches lead to overly broad segmentation masks that include irrelevant parts of the image (i.e., over-segmentation, see Fig. 2 (b)). To alleviate this, [63] employs gradient information from the image-text matching loss, GradCAM[57]-style, to sharpen the attention-based masks for the purpose of guiding caption generation. However, this results in masks that capture only the most discriminative regions of an object, such as the head of an elephant (i.e., under-segmentation, Fig. 2 (c)). To acquire complete object masks, we propose Salience DropOut, which iteratively drops the image patches with high GradCAM attention scores, forcing the model to attend to less discriminative but relevant object parts.
((a))
((b))
((c))
((d))
((a))
((b))
((c))
((d))
Figure 2: Segmentation masks for elephant and train using (a) off-the-shelf cross-attention, (b) cross-attention + GradCAM, and (c) cross-attention + GradCam + Salience DropOut (§3.2). The naive cross-attention masks are too inclusive whereas GradCAM is too exclusive.
Another important consideration is the cross-attention layer and the attention head to extract the masks from. These hyperparameters have enormous influence on the final results and are traditionally tuned on a validation set with pixel-level mask annotations. To eliminate the need for dense annotations, we propose an weakly-supervised reward function based on CLIP [49]. On a validation set with images as well as object class names, the technique contrasts the extracted object regions with a blank image. If, according to CLIP, the former is more similar to the corresponding class name than the latter, we increment the reward. All hyperparameter tuning of our technique is performed with a simple random search with this reward, leading to high performance.
In summary, we propose Plug-and-Play Open-vocabulary Semantic Segmentation (PnP-OVSS), an extremely simple and training-free framework to extract semantic segmentations from VLMs. At zero extra training cost, PnP-OVSS can be used with any pretrained VLM with text-to-image cross attention layer and an image-text matching loss. It has zero reliance on pixel-level annotations, including a validation set for hyperparameter tuning. At the same time, PnP-OVSS delivers excellent performance. It not only beats the training-free baseline with remarkable margins (+29.4% mIoU on Pascal VOC, +13.2% on Pascal Context, +14.0% on MS COCO, +2.4% on COCO Stuff), but also outperforms most recent techniques within the past two years that require extensive finetuning on top of the VLM pretraining.
With this paper, we make three contributions:
  • We propose to combine text-to-image attention, GradCAM, and Salience DropOut to iteratively acquire accurate segmentation of arbitrary classes from pretrained VLM.
  • We replace the densely annotated validation set for hyperparameter tuning, which is needed by most existing methods, with a contrastive reward function based on CLIP. This reward function, coupled with random search, finds a good set of hyperparameters for OVSS.
  • The proposed method, PnP-OVSS, is simple to use, requires no extra finetuning, and delivers high performance. Its success hints at a new direction for open-vocabulary segmentation tasks leveraging large VLMs.
2 Related Work
2.1 Large-Scale Vision-Language Model
Large-scale vision-language models (VLMs), trained on millions of image-text pairs, have become the foundation for many multimodal tasks. Architecturally, the straightforward approach to training such methods involves aligning visual and textual latent representations via a simple dot product [49, 24]. However, this is insufficient for complex structured tasks like visual question answering or image captioning, which require specialized approaches that employ separate encoders before cross-attention between modalities [35, 37, 79, 33, 80] or self-attention networks over all tokens from both modalities [27, 40, 61, 39, 10]. Another design dimension when training VLMs is the loss function. Commonly used losses include image-text contrastive learning [50, 35, 37, 34, 82, 61], image-text matching (ITM) [35, 37, 79, 34, 82, 61, 27, 73, 83], prediction of masked tokens or patches [79, 34, 82, 61], and language modeling [7, 34, 1, 37].
This work utilizes models with unimodal encoders followed by cross-attention fusion [37, 79, 82, 34, 35, 73, 83], as they work with high-level features, and accurately attend to the appropriate image patches. Additionally, we utilize the gradient from the ITM loss in the GradCAM step (§3.1) to sharpen segmentation masks.
2.2 Zero-shot Semantic Segmentation
Method PT Data Size FT data size
Requires finetuning on image-text pairs
OVSegmentor [75] - 4.3M
Vil-Seg [43] 400M 412M
GroupVit* [74] - 3.4M
GroupVit [74] - 26M
CLIPpy [51] - 134M
SegClip [44] 400M 3.4M
ViewCo [52] - 26M
TCL [6] 400M 15M
PACL [46] 400M 30M
Requires finetuning but not image-text pairs
MaskClip w/ ST [90] 400M 1.2M
ZeroSeg* [8] 400M 3.4M
ZeroSeg [8] 400M 1.2M
Requires no finetuning
MaskClip[90] 400M 0
Reco[58] 400M 0
PnP-OVSS (Ours)
+ BLIP 129M 0
+ BridgeTower 400M+ 4M 0
Table 1: Training data of current Zero-shot semantic segmentation methods with only text supervision. PT stands for pretraining with image-caption data, FT stands for fine-tuning with image-caption data. ST stands for self training. We list only the image-caption data used for pretraining and all type of data for finetuning. For hyperparameter tuning, our method uses CLIP-L/14 pretrained with 400M data to calculate the reward.
Zero-shot semantic segmentation predicts a dense segmentation mask for any object class described by a given text prompt, with only prior exposure to class-agnostic image-level supervision. This contrasts with weakly supervised semantic segmentation [88, 3, 13, 54, 28, 62, 84, 26, 70, 23, 70, 30, 87], which relies on class-specific annotations, and unsupervised object discovery [59, 14, 68, 69, 65, 12, 56, 64], which identifies the sole object in the foreground.
Traditional methods [72, 4, 18, 47] train a classifier to distinguish between seen and unseen visual features, wherein the unseen visual features are obtained from a generative model, trained on pairs of seen class image-text embeddings. Recently, methods additionally leverage knowledge from VLMs to attain better matching of visual and textual features [17, 77, 32, 90, 41]. Concretely, they train the segmentation network with dense annotations, while replacing a part of the framework with components from VLMs.
Recently, methods have explored the use of additional supervision to further reduce the need for pixel level annotations. Prior work is compared in Table 1.
Models Finetuned on Image-Text Pairs. Methods proposed in [74, 52, 44, 75, 43, 51, 46, 6] require paired image-text annotations to adapt to the zero-shot segmentation task. Specifically, approaches in [74, 52, 44, 75, 43, 51] cluster semantically similar pixels by contrasting grouped region embeddings with textual embeddings, usually via the use of contrastive and self-supervised losses. PACL [46] modifies the constrastive loss to operate over aggregated patch embeddings (instead of image embeddings) to encourage better alignment between image patches and text. TCL [6] achieves a similar patch-text alignment by introducing an additional module to extract text-grounded image regions.
Models Finetuned on Pseudo-labels. Methods in [8, 90] do not require additional image-text supervision, but involve additional fine-tuning to adapt to the segmentation task. ZeroSeg [8] attempts to match its own group embeddings with multi-scaled segment embeddings obtained from CLIP [49]. MaskClip w/ ST [90] modifies the global attention pooling layer within CLIP to output segmentation masks, which are used as pseudo-labels to train a segmentation network.
In comparison, our proposed PnP-OVSS does not require any fine-tuning or additional paired image-text annotations. It can directly distill high quality open vocabulary semantic segmentations from any VLM with direct text-to-image cross-attention and an image-text matching loss [37, 79, 82, 34, 35, 73, 83]. Compared to approaches in [58, 90] that perform zero-shot open vocabulary semantic segmentation under a similar no training and no additional annotations paradigm, PnP-OVSS achieves considerably superior performance.
Figure 3: The first iteration with cross-attention + GradCAM + Salience DropOut. The text prompt contains $K$ class names and the image contains $P\times P$ patches. From a cross-attention layer and an attention head in the pretrained VLM, we obtain $K$ attention score maps of size $P\times P$, which are sharpened by GradCAM using gradients from the image-text-matching (ITM) loss. To get more complete predictions, we perform Salience Dropout, which repeatedly zero out image patches of the highest average scores and feeds the remaining patches to the image encoder again, forcing the model to attend to other less discriminative patches. We show example salience maps from all iterations in Fig. 4.
3 Method
Figure 4: An illustration of Salience DropOut, showing GradCAM salience values after each iteration. Black squares in the images indicate dropped patches. We obtain the final result by summing the salience maps from all iterations and applying thresholding, Gaussian blur, and Dense CRF.
PnP-OVSS has four major steps. First, we extract a cross-attention salience map per object class from a VLM. Second, we sharpen the salience map by weighing it with the ITM gradient in the style of GradCAM. Third, we apply Salience DropOut that iteratively completes the salience maps. Fourth, we apply Dense CRF [29] for fine-grained adjustment. The process is illustrated in Fig. 3 and Fig. 4.
In the first step, we feed an image and a text prompt to the pretrained VLM and extract cross-attention maps. The text prompt is “A picture of [class 1] [class 2][class K]”, which includes all $K$ class names of interest in the dataset. The image is divided into $P\times P$ patches. The text prompt and the image first go through the modality-specific encoders respectively, followed by a cross-attention fusion module. In the cross-attention layers, the text encodings serve as query vectors and the image patche encodings are as key and value vectors.
We extract an attention map for each text token to the image patches from a particular cross-attention layer and an attention head. Note that different layers and heads lead to drastic performance differences, and the choices are hyperparameters, tuned using the procedure in §3.4. We exclude attention maps for the first three tokens “A”, “picture”, “of”, which do not describe semantic classes. If a class name contains two or more tokens, we take the mean attention map. This procedure yields an attention tensor of size $K\times P\times P$.
With the correct layer and attention head, we observe that this attention map, when normalized by softmax along the first dimension $K$, can provide passable semantic segmentations. However, it tends to include many patches unrelated to the class being segmented, leading to over-segmentation. Hence, we introduce two refinement steps, GradCAM and Salience DropOut, explained in §3.1 and §3.2. After these steps, we acquire aggregate salience maps for every class. Finally, we conduct local polish to the resultant salience maps using Dense CRF, as described in §3.3.
3.1 Map Sharpening with GradCAM
The off-the-shelf attention maps tend to also cover many patches unrelated to the class name (See Fig. 2 (b)). Prior works [35, 63] further sharpen the attention maps and focus them on class-discriminative regions using a variant of GradCAM [57], which is originally proposed for convolutional networks, but applicable to attention maps. It is worth noting that [35, 63] use the technique for purposes other than semantic segmentation.
The GradCAM method requires a gradient. Here we leverage the image-text matching (ITM) loss, which trains the VLM to classify if an image-text pair match each other or not. Computing the ITM loss requires a label. We use “matching” (as opposed to “not matching”) as the label and compute the gradient of the loss with respect to the attention score. This is equivalent to asking: which attention scores contribute the most to the decision that the image-text pair is matching? Formally, we denote a $P\times P$ attention map for class $k$ as $M^{(k)}$ and the ITM loss as $\mathcal{L}_{\text{ITM}}$. The GradCAM class salience map is computed as
$\widetilde{M}^{(k)}=\max\left(0,\frac{\partial\mathcal{L}_{\text{ITM}}}{\partial M^{(k)}}\right)\otimes M^{(k)},$ (1)
where $\otimes$ denote the component-wise multiplication and $\max(\cdot)$ is also applied component-wise.
3.2 Salience DropOut
As illustrated in Fig. 2 (c), segmentations generated by the GradCAM-style re-weighting of cross-attention are often narrowly focused on the most discriminative regions for a given class. However, the less discriminative regions are still important for the completeness of masks. To compel the VLM to attend to these regions, we propose an iterative technique called Salience DropOut.
Since our task is zero-shot and open-vocabulary, we do not have prior knowledge of the classes present in the image. Hence, we sum up the salience maps $\widetilde{M}^{(k)}$ over all classes $k$, yielding a class-agnostic salience map $U^{(t)}$ for the $t^{\text{th}}$ Salience DropOut iteration, $U^{(t)}=\sum^{K}_{k=1}\widetilde{M}^{(k,t)},$ where $\widetilde{M}^{(k,t)}$ is the GradCAM salience map for class $k$ after the $t^{\text{th}}$ iteration. Next, we zero out the 50% of image patches with the highest values in $U^{(t)}$. On the remaining image, we compute the GradCAM salience maps $\widetilde{M}^{(k,t+1)}$ and their sum, $U^{(t+1)}$. Any image patch previously zeroed out will always receive zero salience in later iterations.
Formally, the set of remaining image patches $S^{(t)}\subseteq\{1,\dots,P\}^{2}$ after the $t^{\text{th}}$ dropout iteration is defined as
$S^{(t)}$ $=S^{(t-1)}\setminus\{(i,j)~{}|~{}U^{(t)}_{ij}\geq\eta\},$ (2)
$\eta$ $=\texttt{median}\left(\{U^{(t)}_{ij}~{}|~{}(i,j)\in S^{(t-1)}\}\right),$ (3)
where $U^{(t)}_{i,j}$ denotes the aggregate salience value of the image patch at row $i$ and column $j$. Additionally, note that the input to the first dropout iteration $S^{0}=\{1,\dots,P\}^{2}$ is the set of all image patches.
We stop at four rounds of dropout, as almost all (93.75%) patches are removed beyond that point. The final output for each class $k$ is the sum over all salience maps across the four dropout iterations, $\hat{M}^{(k)}=\sum_{t=1}^{4}\widetilde{M}^{(k,t)}$.
3.3 Gaussian Blur and Dense CRF
The salience dropout procedure generates $k$ continuous-valued salience maps, one per object class. To filter out small random noise in the salience values, we subsequent apply a straightforward thresholding operation at a predefined value $T$ on the salience values and obtain binary segmentation masks. Nonetheless, the hard thresholding creates jagged segmentations with sharp edges that often do not coincide with object boundaries. One common strategy in zero-shot segmentation is to apply Dense Conditional Random Field (CRF) [29], which makes fine-grained adjustments to the estimated masks by enforcing consistency between nearby image pixels with similar colors.
However, we find that the hard 0/1 labels in the binary masks do not work well as pixel unary potentials for Dense CRF. Hence, we smooth them using a Gaussian kernel with a preset variance $\sigma$, which results in a better initialization for unary terms and accounts for uncertainty of exact segmentation boundary along the patch boundaries.
3.4 Hyperparameter Tuning
Three hyperparameters in PnP-OVSS have the strongest influence on the result, the cross-attention layer $L$, the attention head $H$, and the binary threshold $T$. Traditionally, tuning these hyperparameters requires a validation set with pixel-level labels. However, since our goal is to perform zero-shot open-vocabulary semantic segmentation, this requirement could potentially limit the applicability of the technique. Instead, we propose a weakly supervised reward function for hyperparameter tuning, which only requires a set of images and the class names appearing in each image.
The reward for an image $I$ is calculated as follows. We start with a set of classes present in the image, denoted as $\mathcal{K}(I)$. For each class $k\in\mathcal{K}(I)$, we obtain a segmentation mask $M^{(k)}$ (which can be the GradCAM mask, the Salience DropOut mask, or the Dense CRF mask). Next, we apply the mask to the image $I$ and input the extracted regions $M^{(k)}\otimes I$ into a pretrained neural network $f$, which takes an image and a textual class name as input and produces a similarity score. We calculate the normalized probability that the masked image $M^{(k)}\otimes I$ belongs to the ground-truth class $k$ and contrast with a completely black image $\bm{0}$.
Reward $=\sum_{k\in\mathcal{K}(I)}\mathds{1}[Pr(M^{(k)}\otimes I,k)>Pr(\bm{0},k)],$ (4)
$Pr(I,k)$ $=\frac{\exp(f(I,k))}{\sum_{k^{\prime}\in\mathcal{K}(I)}\exp(f(I,k^{\prime}))},$ (5)
where $\mathds{1}(\cdot)$ is the indicator function. Intuitively, a reward of $1$ is assigned if and only if the image features pulled with the estimated mask for class $k$ bears higher similarity to the class name of $k$ than a black image (which can be interpreted as the prior probability of class $k$).
We sum up the reward for all validation images as the total reward. The best hyperparameters, including the cross-attention layer, the attention head, the threshold $T$, and the variance in the Gausian blur kernel, are determined using a simple random search.
4 Experiments
4.1 Datasets and Implementation Details
Following the previous work for zero-shot semantic segmentation, we adopt validation sets of Pascal-VOC 2012 [16], Pascal Context [45], COCO Object [42], COCO Stuff [5], and ADE20K [89] that contain, respectively, 20 object classes, 59 object and stuff classes, 80 object classes, 171 object and stuff classes, and 150 object and stuff classes to evaluate our framework. To verify its versatility, we apply PnP-OVSSto two high-performance VLMs, BLIP [37] and BridgeTower [79], which have the ITM loss and text-to-image cross-attention. More details are in the supplementary material.
Hyperparameters Start End Step Solution
BLIP
Layer 1 12 1 8
Head 1 12 1 10
Attention Threshold 0.05 0.5 0.1 0.15
BridgeTower
Layer 1 6 1 2
Head 1 16 1 8
Attention Threshold 0.05 0.5 0.1 0.15
Gaussian Blur
Standard Deviation 0.01 0.11 0.02 0.05
Table 2: Search space for hyperparameters
Method Finetuning VLMs HT on Dense Labels Short-side Resolution Pascal VOC-20 Pascal Context-59 COCO Object-80 COCO Stuff-171 ADE 20K-150
Group 1: Methods that require weakly supervised finetuning on image-text data
ViL-Seg [43] - 37.3 18.9 - 18.0 -
CLIPpy [51] 224 52.2 - 32.0 25.5 13.5
SegClip [44] 224 52.6 24.7 26.5 - -
GroupVit (by [51]) 224 28.1 14.8 12.9 - 6.2
GroupVit (by [6]) 448 50.4 18.7 27.5 15.3 9.2
GroupVit [74] 448 52.3 22.4 24.3 - -
ViewCo [52] 448 52.4 23.0 23.5 - -
OVSegmentor [75] 448 53.8 20.4 25.1 - -
TCL [6] +PAMR [2] 448 55.0 30.4 31.6 22.4 17.1
PACL [46] 224$\times$4 72.3 50.1 - 38.8 31.4
Group 2: Methods that require finetuning but not real image-text data
MaskClip w/ ST [90] 336 - 31.1 - 18.0 -
MaskClip w/ ST (by [6]) 448 38.8 23.6 20.6 16.4 9.8
ZeroSeg [8] 448 40.8 20.4 20.2 - -
Group 3: Methods that require no finetuning
MaskClip (by [51]) $\times$ $\times$ 224 22.1 - 13.8 8.1 6.8
MaskClip [90] $\times$ $\times$ 336 - 25.5 - 14.6 -
Reco [58] $\times$ $\times$ 320 - 27.2 - 27.2 -
Reco (by [6]) $\times$ $\times$ 448 25.1 19.9 15.7 14.8 11.2
PnP-OVSS with different VLMs
[…]
BridgeTower $\times$ $\times$ 322 36.4 32.3 24.2 27.6 18.6
BridgeTower $\times$ $\times$ 336 35.3 32.4 24.2 27.6 18.0
BridgeTower $\times$ $\times$ 770 35.2 32.4 24.5 28.0 19.0
Table 3: Zero-shot semantic segmentation performance in mIoU. Group 3 contains the most similar baselines that serve as fair comparisons to PnP-OVSS. Groups 1 and 2 benefit from additional training, extra image-text data, and hyperparameter tuning on dense labels. We use the word “by” followed by a paper citation to indicate results of the same technique reported by different papers.  CLIPpy tests on 133 categories of COCO Stuff while we test all 171 classes of COCO Stuff. ViL-Seg is tested on subset of classes on the three datasets, as detailed in the supplementary.
Method Dense Labels HT on Dense Labels Pascal VOC-20 Pascal Context-59 COCO Stuff-171
[…]
SPNet+ST [72] 25.8 - 26.9
ZS3Net+ST [4] 21.2 20.7 10.6
CaGNet+ST [18] 30.3 - 13.4
STRICT [47] 35.6 - 30.3
LSeg [32] 41.0 - -
SimBase [77] 72.5 - 36.3
MaskCLIP+ w/ ST [90] 86.1 66.7 54.7
Table 4: Comparison of zero-shot semantic segmentation performance on unseen categories with methods trained with dense annotation.
Hyperparameter tuning. We use CLIP VIT-L/14 to calculate the reward. The input resolution is 336$\times$336. The search spaces of the random search and results are shown in Tab. 2. For computation efficiency, we tune the layer, the head, and the threshold with GradCAM masks. With the first three hyperparameters fixed, we tune the Gaussian variance on masks before Dense CRF. We directly adopt Dense CRF hyperparameters from CutLER [67] without tuning.
Baselines. We adopt recent papers on zero-shot open-vocabulary semantic segmentation as baselines. The only baselines strictly comparable to our work are MaskClip [90] and Reco [58], which do not perform any finetuning on pretrained VLMs and do not perform hyperparameter tuning on dense annotations. We call these Group 3. To further expand our scope, we also include two other groups of baselines. Group 2 finetunes VLMs but does not require image-text pairs, including MaskClip with self-training (ST) [90] and ZeroSeg [8]. We include ZeroSeg, trained with ImageNet1K, and ZeroSeg*, trained with CC3M+COCO. Group 1 contains baselines that require training on image-text pairs. For completeness, we also compare against supervised techniques since 2019. To maintain the zero-shot setting, we test them on classes not observed during training. For more details, we refer readers to the the supplementary material.
Ablated Model Pascal Context COCO Stuff
[…]
+ GradCam 21.6 17.5
       + Drop 1 25.1 19.8
       + Drop 2 26.5 20.6
       + Drop 3 27.0 20.9
       + Drop 4 27.2 20.9
       + Drop 4 + Blur 36.8 28.6
       + Drop 4 + Dense CRF 35.3 31.8
       + Drop 4 + Blur + Dense CRF 42.8 32.5
[…]
4.2 Main Results
We show the main results in Tab. 3. As input resolution may influence results and cause unfair comparisons, we label the resolution used by each method. PACL [46] uses an 224 resolution but changes the stride for image patchification from 16 to 4, hence introducing overlapping patches.
PnP-OVSS exhibits excellent performance. Comparing with MaskClip [90] and Reco [58], the two methods that require no additional training and ground truth for hyperparameter tuning, on an equal-resolution basis, we attain +29.4% mIoU on Pascal VOC, +13.2% mIoU on Pascal Context, +14.0% mIoU on COCO Object, and +11.4% mIoU on ADE-20K. Further, PnP-OVSS surpasses all baselines in Group 2. On an equal-resolution basis, we achieve +13.7% mIoU on Pascal Voc, +9.6% on Pascal Context, +9.5% on COCO Object, +11.6% on COCO Stuff, and +12.8% on ADE-20K.
[…]
When applied to BridgeTower [79], PnP-OVSS still surpasses all methods in Group 3 by 10.2% on Pascal VOC, 5.1% on Pascal Context, 8.5% on COCO Object, 0.4% on COCO Stuff, and 6.8% on ADE-20K. This showcases the plug-and-play ability of PnP-OVSS, which excels with different base networks.
[…]
4.3 Ablation Study
[…]
Layer Mean 1 2 3 4 5 6
mIoU 11.6 12.3 11.3 11.0 11.6 12.9
Layer Mean 7 8 9 10 11 12
mIoU 13.7 25.3 25.8 23.4 12.5 5.8
Head in Layer 8 1 2 3 4 5 6
mIoU 16.6 3.8 8.6 20.0 19.8 5.2
Head in Layer 8 7 8 9 10 11 12
mIoU 11.1 6.0 3.8 29.6 17.3 8.0
[…]
4.4 Hyperparameter Sensitivity
The choice of hyperparameters often exerts substantial influence on segmentation performance. Here we quantitatively examine how the choice of cross-attention layers and attention heads may change the segmentation mIoU on COCO Stuff. Tab. 6 shows the results obtained from the average cross-attention maps over all heads in each layer and those from different attention heads.
We make the following observations. First, different layers and heads have drastic performance differences. The best-to-worst difference among all layers is 20, and that among heads in Layer 8 is 23.8, underscoring the importance of hyperparameter tuning. Second, the random search using the proposed reward function correctly identifies the best layer-head combination, even though Layer 8 is not the best layer based on average head performance. This indicates the effectiveness of our method.
5 Conclusions
We propose PnP-OVSS, which extracts the ability of semantic segmentation from opaque VLMs. PnP-OVSS is simple to use, requires no extra finetuning, and delivers high performance, exceeding not only all baselines that requires no finetuning, but also all baselines that do not use image-text pairs in finetuning. Its success hints at a new direction for open-vocabulary segmentation tasks leveraging large VLMs.
6 Acknowledgments
This work has been supported by the Nanyang Associate Professorship and the National Research Foundation Fellowship (NRF- NRFF13-2021-0006), Singapore, and by the Vector Institute for AI, Canada CIFAR AI Chair, and NSERC Discovery Grant. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of the funding agencies.
Supplementary Material
[…]
7 Additional Qualitative Results
We provide additional qualitative results of images from Pascal Context and COCO Stuff in Fig 6. We note that, in the third image of the third row, PnP-OVSS correctly recognizes four loaves of bread whereas the ground truth only annotates one.
The rightmost column of the bottom three rows, separated from the rest by a red dash line, are examples of failure cases. The example on top contains multiple small instances of the same class (people and surfboard), which are understandably hard. The middle example contains multiple instances of people, which causes difficulties for PnP-OVSS to cover all objects. In addition, images with a clutter of different objects and complex texture (as in the last image) often cause a drop in performance.
Figure 6: Qualitative Results of PnP-OVSS + BLIP. Images are from Pascal Context and COCO stuff. The bottom rows show the ground-truth (GT); the rest are our results. The last column of the last three rows, after the red dash line, shows failure cases.
8 Qualitative Results in the Wild
[…]
In particular, we would like to point out the difficulty involved in segmenting the Steamboat Willie image, which is in greyscale and has little texture, depriving the network the ability to use color and texture features. Despite that, PnP-OVSS is able to extract the main components of Mickey Mouse, the Helm and the Deck.
9 PnP-OVSS Implementation Detail
[…]
BridgeTower. We use the ITM branch of BridgeTower_$\_{large}$, which adopts VIT-L/14 as the image encoder, RoBERTa_$\_{large}$ as the text encoder and an 6-layer cross attention encoder. For each cross attention layer of the cross-modal encoder, the hidden size is set to 1,024, and the number of heads is set to 16. We interpolate the positional embedding to allow input resolution of 770 x 770. The model weights are from the bridgetowerlarge-itm-mlm-itc checkpoint from Huggingface.
Random Search. We adopt the random search routine from the Gradient-Free-Optimizers library [^1] https://github.com/SimonBlanke/Gradient-Free-Optimizers [60] with our reward metric (§3.4). To parallelize the search process, we divide the search space into three groups and place each group on a GPU card. We perform 34 search iterations in each group. The best hyperparameter set from the three groups with the highest reward is taken as the final search result.
Class Split for Densely Supervised Models. We follow the most common setting [4, 18, 47, 32, 77, 90] which save pottedplant, sheep, sofa, train, tvmonitor as the 5 unseen classes for Pascal VOC; cow, motorbike, sofa, cat, boat, fence, bird, tvmonitor, keyboard, aeroplane as 10 unseen classes for Pascal Context; frisbee, skateboard, cardboard, carrot, scissors, suitcase, giraffe, cow, road, wallconcrete, tree, grass, river, clouds, playingfield, as 15 unseen classes for COCO Stuff.
10 Inference Speed
MaskClip, Reco, and PnP-OVSS take 0.05s, 4.41s, and 2.46s on average, respectively, for inference on a $320\times 320$ image. We calculate the inference speeds of the three models on a single A6000 GPU with 48GB of RAM. The results are the average of 20 independent runs. Hence, PnP-OVSS achieves substantially better performance without significant increase in inference time.
11 Vil-Seg Evaluation Detail
In Tab 3, different from other methods that require weakly supervised finetuning on image-text data, Vil-Seg is evaluated on subset of datasets. Specifically, the author evaluate their method on 5 classes (potted plant, sheep, sofa, train, tv-monitor) out of the 20 object categories in PASCAL VOC; 4 classes (cow, motorbike, sofa, cat) out of the 59 object categories in PASCAL Context; and 15 classes (frisbee, skateboard, cardboard, carrot, scissors, suitcase, giraffe, cow, road, wall concrete, tree, grass, river, clouds, playing-field) out of the 171 object categories in COCO Stuff dataset.
12 Details of Zero-shot Semantic Segmentation Techniques
We summarize current methods for zero-shot semantic segmentation in Tab 7 to enable straightforward comparison between methods. Specifically, we include the supervision used, whether the method require pretraining and finetuning, the pretraining weight, finetuning data and total data size used in each method.
13 PnP-OVSS with ALBEF and mPLUG
As shown in Tab 8, we further apply PnP-OVSS on two other vision language models with cross-attention and image text matching loss, ALBEF [35] and mPLUG [34]. Nevertheless, the performances are not as good as BLIP or BridgeTower. ALBEF [35] is pretrained with only 14M data with ViT-B whereas BLIP and BridgeTower are pretrained with 129M/404M data with ViT-L. We speculate that vision language models require sufficient numbers of parameters and pretraining data to acquire localization capability. mPLUG [34] is another vision language model pretrained with 14M data and ViT-L. However, mPLUG is trained for both image task and video task. With a relatively smaller amount of pretraining data than BLIP or BridgeTower, as well as image and video dual-modality objectives, mPLUG also does not perform well in image object localization.
Method Supervision Training PT weight Finetuning data PT/T data size
Methods that require finetuning on dense annotations w/o VL Models
SPNet [72] Pixel+Self PT+FT ImageNet Pascal Voc/COCO stuff 119K
ZS3Net [4] Pixel+Self PT+FT - Pascal Voc/ Pascal Context 5K
CaGNet [18] Pixel+Self PT+FT - Pascal Voc/Pascal Context /COCO stuff 123K
STRICT [47] Pixel+Self PT+FT ImageNet Pascal Voc/COCO stuff 119K
Methods that require finetuning on dense annotations w/ VL Models
SimBase [77] Pixel+Self PT+FT Maskformer/FCN + CLIP COCO stuff 400.1M
LSeg [32] Pixel+Self PT+FT ImageNet+CLIP Pascal VOC/COCO/FSS 401.3M
MaskCLIP+ [90] Pixel+Self PT+FT CLIP COCO stuff 400.1M
Methods that require finetuning on image-text pairs
OVSegmentor [75] Text+Self PT+FT DINO+BERT CC4M 4.3M
Vil-Seg [43] Text+Self PT+FT CLIP CC12M 412M
GroupVit* [44] Text T - CC3M+COCO 3.4M
GroupVit [74] Text T - CC12M+YFCC14M 26M
CLIPpy [51] Text+Text PT+FT DINO+T5 Sentence HQITP-134M 134M
SegCLIP [44] Text+Self PT+FT CLIP CC3M+COCO 403.4M
Viewco [52] Text PT+FT GroupViT CC12M+YFCC14M 26M
TCL[6]+PAMR[2] Text PT+FT CLIP CC12M+CC3M 415M
PACL [46] Text PT+FT CLIP GCC3M+GCC12M +YFCC15M 430M
Methods that require finetuning but not image-text pair
MaskCLIP [90] w/ ST Text+ST PT+T CLIP ImageNet1K 401.2M
ZeroSeg* [8] Text+Self PT+T CLIP CC3M+COCO 403.4M
ZeroSeg [8] Text+Self PT+T CLIP ImageNet1K 401.2M
Methods that require no finetuning
MaskCLIP [90] Text PT CLIP - 400M
Reco[58] Text PT CLIP+ImageNet - 400M
PnP-OVSS (Ours)
[…]
     + BridgeTower Text PT BridgeTower - 404M
Table 7: Current methods for zero-shot semantic segmentation. Pixel represents method require pixel level annotation, Self represents method leverage self supervision, and Text represents method leverage image-text pair annotation. PT stands for Pre-training, T stands for training, ST stands for Self-training, FT stands for finetuning. All methods with pixel supervision are trained on seen categories and tested on unseen categories. For data size, We calculate only the image-caption data used for pretraining and all type of data for finetuning.
Method Training HT on Dense Labels Short-side Resolution Pascal VOC-20 Pascal Context-59 COCO Object-80 COCO Stuff-171
PnP-OVSS (Ours)
     + ALBEF $\times$ $\times$ 336 10.8 6.3 8.9 10.3
     + ALBEF $\times$ $\times$ 768 11.1 6.8 8.8 10.7
     + mPLUG $\times$ $\times$ 336 9.2 8.8 7.9 6.7
Table 8: Zero-shot semantic segmentation performance in mIoU.

### v3

@@
 PnP-OVSS does not require any neural network training and performs hyperparameter tuning without the need for any segmentation annotations, even for a validation set.
-PnP-OVSS demonstrates substantial improvements over comparable baselines (+29.4% mIoU on Pascal VOC, +13.2% mIoU on Pascal Context, +14.0% mIoU on MS COCO, +2.4% mIoU on COCO Stuff) and even outperforms most baselines that conduct additional network training on top of pretrained VLMs.
+PnP-OVSS demonstrates substantial improvements over comparable baselines (+26.2% mIoU on Pascal VOC, +20.5% mIoU on MS COCO, +3.1% mIoU on COCO Stuff and +3.0% mIoU on ADE20K).
 Our codebase is at https://github.com/letitiabanana/PnP-OVSS.
 Figure 1: Qualitative Results of PnP-OVSS + BLIP.
-Images are from Pascal Context and COCO stuff.
+Images are from Pascal VOC and COCO Object.
 The right columns and bottom rows show the ground-truth (GT); the rest are our results.
-Note accurate results even on complex (trees) and small objects (last column).
+Note the good results on small objects like the frisbee and the tennis racket.
 1 Introduction
@@
 ((d))
-Figure 2: Segmentation masks for elephant and train using (a) off-the-shelf cross-attention, (b) cross-attention + GradCAM, and (c) cross-attention + GradCam + Salience DropOut (§3.2).
+Figure 2: Segmentation masks for elephant and giraffe using (a) off-the-shelf cross-attention, (b) cross-attention + GradCAM, and (c) cross-attention + GradCam + Salience DropOut (§3.2).
 The naive cross-attention masks are too inclusive whereas GradCAM is too exclusive.
@@
 At the same time, PnP-OVSS delivers excellent performance.
-It not only beats the training-free baseline with remarkable margins (+29.4% mIoU on Pascal VOC, +13.2% on Pascal Context, +14.0% on MS COCO, +2.4% on COCO Stuff), but also outperforms most recent techniques within the past two years that require extensive finetuning on top of the VLM pretraining.
+It not only beats the training-free baseline with remarkable margins (+26.2% mIoU on Pascal VOC, +20.5% mIoU on MS COCO, +3.1% mIoU on COCO Stuff and +3.0% mIoU on ADE20K), but also outperforms most recent techniques within the past two years that require extensive finetuning on top of the VLM pretraining.
 With this paper, we make three contributions:
@@
 We obtain the final result by summing the salience maps from all iterations and applying thresholding, Gaussian blur, and Dense CRF.
-PnP-OVSS has four major steps.
-First, we extract a cross-attention salience map per object class from a VLM.
-Second, we sharpen the salience map by weighing it with the ITM gradient in the style of GradCAM.
-Third, we apply Salience DropOut that iteratively completes the salience maps.
-Fourth, we apply Dense CRF [29] for fine-grained adjustment.
+PnP-OVSS has five major steps.
+First, we prompt GPT4o to output the categories appear in each image.
+Second, we extract a cross-attention salience map for each predicted category from a VLM.
+Third, we sharpen the salience map by weighing it with the ITM gradient in the style of GradCAM.
+Fourth, we apply Salience DropOut that iteratively completes the salience maps.
+Finally, we apply Dense CRF [29] for fine-grained adjustment.
 The process is illustrated in Fig.
@@
 4.
-In the first step, we feed an image and a text prompt to the pretrained VLM and extract cross-attention maps.
-The text prompt is “A picture of [class 1] [class 2][class K]”, which includes all $K$ class names of interest in the dataset.
+Directly inputting all the labels from the dataset into the VLM could introduce unnecessary noise into the segmentation process.
+In the first step, we prompt GPT4o to filter the possible categories within each image.
+The prompt for GPT4o is: “Help me find as much and accurate as possible, categories appear in the image.
+The the available categories are {CATEGORIES}.
+Double check if there is anything missing.
+You may output unsure categories and give them lower probability.
+Output strictly in the format [id1: classname1, id2: classname2, …], [probability of class1 in percentage, probability of class2 in percentage, …] without other words.
+” {CATEGORIES} being a placeholder for all the categories in a dataset.
+The categories output by GPT4o with higher than 70% probability are treated as the predicted classes for the image.
+In the second step, we feed an image and a text prompt to the pretrained VLM and extract cross-attention maps.
+The text prompt is “A picture of [class 1] [class 2][class K]”, which includes the $K$ GPT4o predicted categories of the image.
 The image is divided into $P\times P$ patches.
@@
 This procedure yields an attention tensor of size $K\times P\times P$.
-With the correct layer and attention head, we observe that this attention map, when normalized by softmax along the first dimension $K$, can provide passable semantic segmentations.
+With the correct layer and attention head, we observe that this attention map can provide passable semantic segmentations.
 However, it tends to include many patches unrelated to the class being segmented, leading to over-segmentation.
@@
 To compel the VLM to attend to these regions, we propose an iterative technique called Salience DropOut.
-Since our task is zero-shot and open-vocabulary, we do not have prior knowledge of the classes present in the image.
-Hence, we sum up the salience maps $\widetilde{M}^{(k)}$ over all classes $k$, yielding a class-agnostic salience map $U^{(t)}$ for the $t^{\text{th}}$ Salience DropOut iteration, $U^{(t)}=\sum^{K}_{k=1}\widetilde{M}^{(k,t)},$ where $\widetilde{M}^{(k,t)}$ is the GradCAM salience map for class $k$ after the $t^{\text{th}}$ iteration.
-Next, we zero out the 50% of image patches with the highest values in $U^{(t)}$.
+We sum up the salience maps $\widetilde{M}^{(k)}$ over the predicted classes $K$, yielding a class-agnostic salience map $U^{(t)}$ for the $t^{\text{th}}$ Salience DropOut iteration, $U^{(t)}=\sum^{K}_{k=1}\widetilde{M}^{(k,t)},$ where $\widetilde{M}^{(k,t)}$ is the GradCAM salience map for class $k$ after the $t^{\text{th}}$ iteration.
+Next, we zero out 10 image patches with the highest values in $U^{(t)}$.
 On the remaining image, we compute the GradCAM salience maps $\widetilde{M}^{(k,t+1)}$ and their sum, $U^{(t+1)}$.
@@
 $\eta$ $=\texttt{median}\left(\{U^{(t)}_{ij}~{}|~{}(i,j)\in S^{(t-1)}\}\right),$ (3)
+$S^{(t)}$ $=S^{(t-1)}\setminus\{(i,j)~{}|~{}U^{(t)}_{ij}\in A\},$ (4)
+[…]
 where $U^{(t)}_{i,j}$ denotes the aggregate salience value of the image patch at row $i$ and column $j$.
 Additionally, note that the input to the first dropout iteration $S^{0}=\{1,\dots,P\}^{2}$ is the set of all image patches.
-We stop at four rounds of dropout, as almost all (93.75%) patches are removed beyond that point.
-The final output for each class $k$ is the sum over all salience maps across the four dropout iterations, $\hat{M}^{(k)}=\sum_{t=1}^{4}\widetilde{M}^{(k,t)}$.
+We stop at the third rounds of dropout. The final output for each class $k$ is the sum over all salience maps across the three dropout iterations, $\hat{M}^{(k)}=\sum_{t=1}^{4}\widetilde{M}^{(k,t)}$.
 3.3 Gaussian Blur and Dense CRF
@@
 ViewCo [52] 448 52.4 23.0 23.5 - -
-OVSegmentor [75] 448 53.8 20.4 25.1 - -
+OVSegmentor [75] 448 53.8 20.4 25.1 - 5.6
 TCL [6] +PAMR [2] 448 55.0 30.4 31.6 22.4 17.1
@@
 MaskClip w/ ST (by [6]) 448 38.8 23.6 20.6 16.4 9.8
+ZeroSeg* [8] 448 37.3 19.7 17.8 -
 ZeroSeg [8] 448 40.8 20.4 20.2 - -
@@
 […]
-[…]
-[…]
-[…]
-[…]
-BridgeTower $\times$ $\times$ 322 36.4 32.3 24.2 27.6 18.6
-BridgeTower $\times$ $\times$ 336 35.3 32.4 24.2 27.6 18.0
-BridgeTower $\times$ $\times$ 770 35.2 32.4 24.5 28.0 19.0
+BridgeTower $\times$ $\times$ 336 42.4 25.3 30.4 15.7 14.8
 Table 3: Zero-shot semantic segmentation performance in mIoU.
@@
 We use the word “by” followed by a paper citation to indicate results of the same technique reported by different papers.
-CLIPpy tests on 133 categories of COCO Stuff while we test all 171 classes of COCO Stuff.
+CLIPpy tests on 133 categories of COCO Stuff and Reco tests on 27 super categories while we test all 171 classes of COCO Stuff.
 ViL-Seg is tested on subset of classes on the three datasets, as detailed in the supplementary.
@@
 […]
-+ GradCam 21.6 17.5
-       + Drop 1 25.1 19.8
-       + Drop 2 26.5 20.6
-       + Drop 3 27.0 20.9
-       + Drop 4 27.2 20.9
-       + Drop 4 + Blur 36.8 28.6
-       + Drop 4 + Dense CRF 35.3 31.8
-       + Drop 4 + Blur + Dense CRF 42.8 32.5
++ GradCam 20.2 14.4
+       + Drop 1 21.6 14.9
+       + Drop 2 22.4 15.2
+       + Drop 3 22.9 15.3
+       + Drop 3 + Blur 26.1 17.1
+       + Drop 3 + Dense CRF 25.1 16.0
+       + Drop 3 + Blur + Dense CRF 28.9 17.9
 […]
@@
 PnP-OVSS exhibits excellent performance.
-Comparing with MaskClip [90] and Reco [58], the two methods that require no additional training and ground truth for hyperparameter tuning, on an equal-resolution basis, we attain +29.4% mIoU on Pascal VOC, +13.2% mIoU on Pascal Context, +14.0% mIoU on COCO Object, and +11.4% mIoU on ADE-20K.
-Further, PnP-OVSS surpasses all baselines in Group 2.
-On an equal-resolution basis, we achieve +13.7% mIoU on Pascal Voc, +9.6% on Pascal Context, +9.5% on COCO Object, +11.6% on COCO Stuff, and +12.8% on ADE-20K.
+Comparing with MaskClip [90] and Reco [58], the two methods that require no additional training and ground truth for hyperparameter tuning, on an equal-resolution basis, we attain +26.2% mIoU on Pascal VOC, +0.8% mIoU on Pascal Context, +20.5% mIoU on COCO Object, +3.1% mIoU COCO Stuff and +3.0% mIoU ADE-20K.
+Further, PnP-OVSS surpasses all baselines except MaskClip with ST in Group 2.
 […]
-When applied to BridgeTower [79], PnP-OVSS still surpasses all methods in Group 3 by 10.2% on Pascal VOC, 5.1% on Pascal Context, 8.5% on COCO Object, 0.4% on COCO Stuff, and 6.8% on ADE-20K.
+When applied to BridgeTower [79], PnP-OVSS still surpasses all methods in Group 3 by 17.3% on Pascal VOC, 14.7% on COCO Object, 0.9% on COCO Stuff, and 3.6% on ADE-20K.
 This showcases the plug-and-play ability of PnP-OVSS, which excels with different base networks.
@@
 Layer Mean 1 2 3 4 5 6
-mIoU 11.6 12.3 11.3 11.0 11.6 12.9
+mIoU 18.6 16.4 12.5 21.2 20.4 19.5
 Layer Mean 7 8 9 10 11 12
-mIoU 13.7 25.3 25.8 23.4 12.5 5.8
+mIoU 21.5 25.2 23.6 23.0 11.7 0.8
 Head in Layer 8 1 2 3 4 5 6
-mIoU 16.6 3.8 8.6 20.0 19.8 5.2
+mIoU 10.4 2.0 3.3 29.8 26.2 6.5
 Head in Layer 8 7 8 9 10 11 12
-mIoU 11.1 6.0 3.8 29.6 17.3 8.0
+mIoU 12.8 3.6 3.0 32.3 12.4 6.2
 […]
@@
 First, different layers and heads have drastic performance differences.
-The best-to-worst difference among all layers is 20, and that among heads in Layer 8 is 23.8, underscoring the importance of hyperparameter tuning.
-Second, the random search using the proposed reward function correctly identifies the best layer-head combination, even though Layer 8 is not the best layer based on average head performance.
-This indicates the effectiveness of our method.
+The best-to-worst difference among all layers is 24.4, and that among heads in Layer 8 is 30.3, underscoring the importance of hyperparameter tuning.
+Second, the random search using the proposed reward function correctly identifies the best layer-head combination This indicates the effectiveness of our method.
 5 Conclusions
 We propose PnP-OVSS, which extracts the ability of semantic segmentation from opaque VLMs.
-PnP-OVSS is simple to use, requires no extra finetuning, and delivers high performance, exceeding not only all baselines that requires no finetuning, but also all baselines that do not use image-text pairs in finetuning.
+PnP-OVSS is simple to use, requires no extra finetuning, and delivers high performance, exceeding not only all baselines that requires no finetuning, but also many of the baselines that require finetuning.
 Its success hints at a new direction for open-vocabulary segmentation tasks leveraging large VLMs.
@@
 We provide additional qualitative results of images from Pascal Context and COCO Stuff in Fig 6.
-We note that, in the third image of the third row, PnP-OVSS correctly recognizes four loaves of bread whereas the ground truth only annotates one.
-The rightmost column of the bottom three rows, separated from the rest by a red dash line, are examples of failure cases.
-The example on top contains multiple small instances of the same class (people and surfboard), which are understandably hard.
-The middle example contains multiple instances of people, which causes difficulties for PnP-OVSS to cover all objects.
-In addition, images with a clutter of different objects and complex texture (as in the last image) often cause a drop in performance.
+We note that, in the third image of the third row, PnP-OVSS correctly recognizes the occluded person missing in the ground truth.
+The bottom row are examples of failure cases.
+The examples on top contains multiple small instances of the same class (people and forks, which are understandably hard.
+All the example contains multiple instances of people, which causes difficulties for PnP-OVSS to cover all objects.
+In addition, images with a clutter of different objects and complex texture often cause a drop in performance.
 Figure 6: Qualitative Results of PnP-OVSS + BLIP.
-Images are from Pascal Context and COCO stuff.
+Images are from Pascal Context, COCO Object and COCO stuff.
 The bottom rows show the ground-truth (GT); the rest are our results.
-The last column of the last three rows, after the red dash line, shows failure cases.
+The last row shows failure cases.
 8 Qualitative Results in the Wild